OpenStudy (gorica):

h(t) is once differentiable function on [0,1] and h(0)=h(1)=0. I'll write question in comment.

OpenStudy (gorica):

Why \[\int\limits_{0}^{1}h'(t)dt=0\] but \[\int\limits_{0}^{1}h'(t)^{2}dt \ge0\] ?

OpenStudy (anonymous):

Well the first one is easy right? \[\int\limits_0^1 \frac{d h(t)}{dt}dt = h(1)-h(0); \text{but } h(1) = 0 =h(0); \text{so} h(1)-h(0)=0\]

OpenStudy (gorica):

ok

OpenStudy (anonymous):

Well we know that: \[0=\left| \int\limits_0^1 h(t)dt\right| \le \left| \int\limits_0^1 (h(t))^2dt\right|\] For any function h(t), so since int of h(t)^2 is positive definite you can drop the magnitude and that implies the integral is greater than or equal to 0.

OpenStudy (anonymous):

I feel as though these are the "qualitative' arguments that I get from my professors when it comes to stuff like this.

OpenStudy (gorica):

thank you :)

OpenStudy (gorica):

I have one more question. Why \[3\int\limits_{0}^{1}h'(t)^{2}(1+\frac{ 1 }{ 3 }h'(t)^{3})dt \ge0\] if \[||h||_{1}\le3\] where \[||h||_{1}=\max|h(t)|+\max|h'(t)|, t \in[0,1]\]